9.48 L of a gas was at an unknown pressure. However, at standard pressure, its volume was measured to be 8.00 L. What was the unknown pressure?
1
Expert's answer
2021-10-20T02:51:28-0400

P1V1= P2V2

Standard pressure= 1atm

1×8= 9.48× P2

8 = 9.48p2

P2= 8/9.48

= 0.844atm